OpenStudy (anonymous):

Simplify -4(x -5) + 3 (a - 7) A.-4x + 3a + 1 B. -4x + 3a + 41 C. -4x + 3a - 41 D. -4x + 3a - 1

alones (alones):

Okay −4(x−5)+3(a−7) first istribute: =(−4)(x)+(−4)(−5)+(3)(a)+(3)(−7) =−4x+20+3a+−21 Combine Like Terms: =−4x+20+3a+−21 =(3a)+(−4x)+(20+−21) =3a+−4x+−1

pooja195 (pooja195):

Don't over think it simply distribute 

alones (alones):

Oh haha xP

OpenStudy (anonymous):

Thanks for the help you both! I appreciate it! :) I understand better now.